You are saving dimes and quarters in your piggy bank. You have 42 coins that have a total value of $6.75. How many dimes and quarters do you have in your piggy bank?
Answer by addingup(3677)

d+q = 42 subtract d from both sides: q = 42-d
and
0.10d+0.25q = 6.75 substitute here for q:
0.10d+0.25(42-d) = 6.75
0.10d+10.50-0.25d = 6.75
-0.15d = -3.75 divide both sides by -0.15 and remember -/- = +
d = 25 So you have 25 dimes and:
42-25 = 17 quarters
---------------------------------
Check:
0.10(25) = 2.50
0.25(17) = 4.25
Total . . 6.75 Correct